OpenStudy (anonymous):

A gallon of water weights 8 1/3 Ib. How much does 1 1/5 gallon weigh? A. 12 Ib B. 8 1/15 Ib C. 10 Ib D. 11 1/2 Ib

OpenStudy (whpalmer4):

what is 1 1/5 gallon * 8 1/3 lbs/gallon? you'll want to convert both numbers to improper fractions before multiplying.

OpenStudy (anonymous):

idk -_-

OpenStudy (anonymous):

I think it's A

OpenStudy (whpalmer4):

oh, come on, this isn't difficult to do. 8 1/3 = (8*3+1)/3 = 25/3 1 1/5 = (1*5+1)/5 = 6/5 25/3 * 6/5 = (25*6)/(3*5) = 150/15 = ???

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

omg I added to the wrong thing lol he is right

OpenStudy (whpalmer4):

yes, 10 is correct.

OpenStudy (whpalmer4):

as you have answer choices here that aren't all that close together, you could just estimate it, too. 1* 8 1/3 = 8 and a bit. add another 1/5 of that, that's about 1.6. 8 and a bit plus almost 2 gives you about 10. B is obviously wrong, it isn't nearly big enough, and A and D are too big.

OpenStudy (anonymous):

Thank you!!!